solve the following system of equations algebraically
3x-1=0
5x+2y=22
part 1: multiply one of the equations by a constant in order to solve by elimination. show the new system of equations below?
part 2: combine the two equations to eliminate one of the variables. show the result go this combination.
part 3: solve for the variable. what is the value?
part 4: substitute this value back into an equation to solve for other variable. write the solution to this system below

Answer by macston(5194)

.
3x-1=0
5x+2y=22
.
part 1: multiply one of the equations by a constant in order to solve by elimination. show the new system of equations below?
Multiply first equation by 5/3:
5x-5/3=0
.
part 2: combine the two equations to eliminate one of the variables. show the result go this combination.
(5x+2y)-(5x-5/3)=22-0
2y+5/3=22
2y=22-5/3
2y=66/3-5/3=61/3
y=61/6
.
part 3: solve for the variable. what is the value?
y=61/6
.
part 4: substitute this value back into an equation to solve for other variable. write the solution to this system below
.
Second equation:
5x+2y=22
5x+(2)(61/6)=22
5x+122/6=22
5x=22-122/6
5x=132/6-122/6=10/6
x=10/30
x=1/3
_______________
CHECK:
.
Solve first equation for x:
3x-1=0
3x=1
x=1/3
.
Substitute in second equation:
5x+2y=22
5/3+2y=22
2y=66/3-5/3
2y=61/3
y=61/6